You have found an asset with an arithmetic average return of 12.80 percent and a geometric average return of 10.28 percent. Your observation period is 40 years. What is your best estimate of the return of the asset over the next 5 years? 10 years? 20 years? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)
Answer #1

Answer:

5 year estimate = [(5 – 1)/(40 – 1)] × 10.28% + [(40 – 5)/(40 – 1)] × 12.80% = 12.54%

10 year estimate = [(10 – 1)/(40 – 1)] × 10.28% + [(40 – 10)/(40 – 1)] × 12.80% = 12.22%

20 year estimate = [(20 – 1)/(40 – 1)] × 10.28% + [(40 – 20)/(40 – 1)] × 12.80% = 11.57%